The distance between object and screen is D.Real images of an object are formed on the screen for two positions of a lens seperated by a distance d.The ratio between the sizes of two images will be-
  1. D/d
  2. D2/d2
  3. (D-d)2/(D+d)2

Grade:12

3 Answers

Mayajyothi C J
7 years ago
 m=\frac{height of image}{height of object}
Height of first image= h1
height of second image=h2
Magnification of first image,
m=\frac{h_{1}}{h_{0}}=\frac{D+d}{D-d}    …...............................(1)
Magnification of second image,
m_{2}=\frac{h_{2}}{h_{0}}=\frac{D-d}{D+d}….....................................(2)
eq (1) divided by eq(2)
\frac{m_{1}}{m_{2}}=\frac{h_{1}}{h_{0}}\div \frac{h_{2}}{h_{0}}
 
Mayajyothi C J
7 years ago
 m=\frac{height of image}{height of object}
Height of first image= h1
height of second image=h2
Magnification of first image,
m=\frac{h_{1}}{h_{0}}=\frac{D+d}{D-d}    …...............................(1)
Magnification of second image,
m_{2}=\frac{h_{2}}{h_{0}}=\frac{D-d}{D+d}….....................................(2)
eq (1) divided by eq(2)
\frac{m_{1}}{m_{2}}=\frac{h_{1}}{h_{0}}\div \frac{h_{2}}{h_{0}}
        =\frac{h_{1}}{h_{2}}= \frac{\frac{D+d}{D-d}}{\frac{D-d}{D+d}}
        =\frac{(D+d)^{2}}{(D-d)^{2}}
The ratio between the size of two image can mean h1/h2 or h2/h1 right? As it is not specified,
Therefore,
  
 
 
Mayajyothi C J
7 years ago
 m=\frac{height of image}{height of object}
Height of first image= h1
height of second image=h2
Magnification of first image,
m=\frac{h_{1}}{h_{0}}=\frac{D+d}{D-d}    …...............................(1)
Magnification of second image,
m_{2}=\frac{h_{2}}{h_{0}}=\frac{D-d}{D+d}….....................................(2)
eq (1) divided by eq(2)
\frac{m_{1}}{m_{2}}=\frac{h_{1}}{h_{0}}\div \frac{h_{2}}{h_{0}}
        =\frac{h_{1}}{h_{2}}= \frac{\frac{D+d}{D-d}}{\frac{D-d}{D+d}}
        =\frac{(D+d)^{2}}{(D-d)^{2}}
The ratio between the size of two image can mean h1/h2 or h2/h1 right? As it is not specified,
Therefore,
  \frac{h_{2}}{h_{1}}=\frac{(D-d)^{2}}{(D+d)^{2}}
Please approve ^_^